• Main INDEX
  • Monthly INDEX
  • PREV
  • NEXT

    User name S. Kowalski

    Log entry time 16:15:18 on July30,2005

    Entry number 150112

    keyword=Shift Summary

    8:00 am Continue production running. Current Run = 3967.
    8:25 Lost Left Spectrometer IOC. Lost Q2, Q3 and Dipole currents.
    Left Septum current reset.
    9:13 Back in production again. Took a while to get everything reset.
    10:38 Rebooted Compton DAQ.
    12:25 ROC 26 crashed. Stopped the run. Does not reboot from Vxworks
    prompt, Bryan tries; call for controlled access to reboot.

    13:00 rebooted ROC26. Arrive back in counting room, find septum
    IOC has lost communication - the voice of a higher authority
    (R.O. Michaels) tells that "the sweeper magnet current is too
    low" - this is because the septum IOC (which also reads out the sweeper magnet). Do a hareboot on the IOC. Communications comes back after the reboot, but the Lakeshore can't read the temperatures
    so we can't find out if the septa have quenched. It says the
    "Lakeshore is being initialized... please wait 15 minutes".
    Afetr waiting patiently, this never clears. So, we click the
    Power Supply ON and then reset, and then everything clears, the
    temperatures are OK, they have not quenched (but currents were
    zero). We ask for the canonical currents and they start
    ramping up).

    13:34 Septa ramping up to final currents. Beam on to check position and size. Spot ++ = 4x2 cm. Beam at 36 micro-A. Unfortunately septum does not cool well at low currents. Exceeded temp limits. Shut off beam immediately.

    13:51 Production run started. Septum seems to be cooling better than earlier. Increased current to 38 micro-A. See what happens.

    15:30 Lost Happex DAQ. Need an access.